OpenStudy (anonymous):

Math help! will reward medal! The system of equations is coincident. What are the missing values? -2x+5y=-12 -6x+____ y=_____

OpenStudy (anonymous):

@K12girlohva please help!

OpenStudy (michele_laino):

please multiply both sides of first equation by \(3\)

OpenStudy (anonymous):

what do you mean? @Michele_Laino

OpenStudy (michele_laino):

hint: \[\huge\left( { - 2x + 5y} \right) \cdot 3 = \left( { - 12} \right) \cdot 3\] please simplify

OpenStudy (anonymous):

I still don't quite understand how to do that exactly...... @Michele_Laino

OpenStudy (michele_laino):

next step: if I apply the distributive property of multiplication ove addition, I get: \[\huge \left( { - 2} \right) \cdot 3x + 5 \cdot 3y = - 36\] please continue

OpenStudy (michele_laino):

what is \((-2) \cdot 3=...?\)

OpenStudy (michele_laino):

what is \(5 \cdot 3=...?\)

OpenStudy (anonymous):

the first one: -6 the second one: 15

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

so we can write this: \(\huge -6x+15y=-36\) so, what are the missing coefficients

OpenStudy (michele_laino):

please compare such equation with the incomplete second equation which you provided

OpenStudy (anonymous):

x and y @Michele_Laino

OpenStudy (mathstudent55):

If the system is coincident, that means the two equations are the same line. That means that one equation is simply a multiple of the other equation. \(−2x+5y=−12\) \(−6x+\_\_\_\_y=\_\_\_\_\) If the second equation is a multiple of the first equation, there is a number that you multiply the first equation by to get the second equation. The only corresponding coefficients that you see in both equations are the coefficients of x: -2 and -6. What do you multiply -2 by to get -6? That's what you need to multiply the other coefficients of the first equation to get the second equation.

OpenStudy (anonymous):

3 @mathstudent55

OpenStudy (phi):

do you get the idea? the 2nd equation has a missing 15 (in front of the y) and -36 on the right side of the = sign.

OpenStudy (mathstudent55):

Correct. Now that you know the number is 3, multiply the other coefficients of the first equation by 3 to get the second equation. \(−2x+5y=-12\) \(3 \times (−2x)+ 3 \times 5 y=3 \times (-12)\) \(-6x + 15y - 36\)